What is the price of a 3 -year, 8.1% coupon rate, $1,000 face value bond that pays interest quarterly if the yield to maturity on similar bonds is 11.9 %

What is the yield to maturity of at 9.9% semiannual coupon bond with a face value of $1,000 selling for $894.71 that matures in 11 years?

The annual yield to maturity of the bond is: (Select the best choice below.)

A.11.62%

B.11.039%

C.12.201%

D.13.131%

E.9.528%
Answer #1

1.Information provided:

Face value= future value= $1,000

Time= 3 years

Yield to maturity= 11.9%

Coupon rate= 8.1%

Coupon payment= 0.081*1,000= $81

The price of the bond is calculated by computing the present value.

Enter the below in a financial calculator to compute the present value:

FV= 1,000

N= 3

I/Y= 11.9

   CPT= 81

Press the CPT key and PV to compute the present value.

The value obtained is 908.57.

Therefore, the price of the bond is $908.57.

2.Information provided:

Face value= future value= $1,000

Current price= present value= $894.71

Coupon rate= 9.9%/2= 4.95%

Coupon payment= 0.0495*1,000= $49.50

Time= 11 years*2= 22 semi-annual periods

The yield to maturity is calculated by entering the below in a financial calculator:

FV= 1,000

PV= -894.71

PMT= 49.50

N= 22

Press the CPT key and I/Y to compute the yield to maturity.

The value obtained is 5.81.

Therefore, the yield to maturity is 5.81%*2= 11.62%.

Hence, the answer is option A.

In case of any query, kindly comment on the solution.